# The evolution and detection of vector superradiant instabilities

### v1

@@ IV Multi-mode evolution
For the $j=1$ modes, the superradiance rate in Eq. (23) is proportional to $(M\mu)^{7}$, while the GW emission rate is proportional to $(M\mu)^{10}$. For the $j=2$ modes, the superradiance rate is $\sim(M\mu)^{11}$, and the emission rate is $\sim(M\mu)^{14}$. In the $M\mu\ll 1$ limit, the $j=2$ modes become crucial at the time when the $j=1$ modes are almost drained by GW emission. As a result, the evolution of these four modes can be separated into two stages. In the first stage, only the $j=1$ modes are important while the $j=2$ modes can be ignored. Then in the second stage, the $j=1$ modes are almost depleted and one only needs to consider the $j=2$ modes. Such hierarchy also leads to different phases in each stage similar to the single-mode evolution. In fact, we will see that the effect of the subdominant mode is so small that the evolutions of the dominant mode and the BH are almost the same as the single-mode case detailed in Section III.
@@ IV.1 The $j=1$ stage
where $\Gamma_{1011}/\Gamma_{0011}\sim\beta_{10}/\beta_{00}=1/8$ with,
@@
It means the mass ratio of the subdominant mode to the dominant mode at $t_{1}$ is smaller if the latter is larger in size. This ratio is important for the modulation of the GW emission in Eq. (30), which will be discussed in Section V.
@@
During this time, the BH mass and spin drop rapidly to $M_{1}$ and $a_{*\text{c}}^{(1011)}$, respectively. Since the subdominant mode is much smaller in mass than the dominant one, one could make estimates of $M_{1}$ and $M_{\text{v}}^{0011}(t_{1})$ as if only the dominant mode exists. Then the problem reduces to the single-mode evolution explained in Section III.1. The $M_{\text{v,max}}^{(0011)}$ and $a_{*\text{c}}^{(0011)}$ in Eqs. (42) should be replaced by $M_{\text{v}}^{(0011)}(t_{1})$ and $a_{*\text{c}}^{(1011)}$, respectively. Nonetheless, using $a_{*\text{c}}^{(1011)}$ or $a_{*\text{c}}^{(0011)}$ here leads to an error only at order $\mathcal{O}\left((M\mu)^{8}\right)$. As a result, Eq. (43) can be used to estimate $M_{\text{v}}^{(0011)}(t_{1})$.
With $M_{\text{v}}^{(0011)}(t_{1})$ calculated, one could further obtain $M_{\text{v}}^{(1011)}(t_{1})$ with Eq. (51) and obtain $t_{1}$ with Eq. (50). Apparently, $M_{\text{v}}^{(1011)}(t_{1})$ is the maximum mass of the subdominant mode. The BH mass at time $t_{1}$ is then,
$M_{1}$ $\approx M_{0}-{M_{\mathrm{v}}^{(1011)}}(t_{1})-{M_{\mathrm{v}}^{(0011)}}(t_{1}),$ (53)
The second phase starts at $t_{1}$ and ends at $t_{2}$ when the subdominant mode is drained. In this phase, the BH spin is between $a_{*\text{c}}^{(1011)}$ and $a_{*\text{c}}^{(0011)}$. The subdominant mode gradually falls into the BH while the dominant mode still extracts energy out of the horizon. At $t_{2}$, the BH spin equals to $a_{*\text{c}}^{(0011)}$. In this phase, the BH mass is almost unchanged, meaning the balance of the inflow and the outflow at the horizon,
@@
Since $M_{\mathrm{v}}^{(0011)}(t)\gg M_{\mathrm{v}}^{(1011)}(t)$, the BH spin in this time range is very close to $a_{*\text{c}}^{(0011)}$. As an approximation, we assume the BH spin is always $a_{*\text{c}}^{(0011)}$ in this time range. Then the superradiance rate $\Gamma_{0011}=0$, indicating the horizon angular speed $\Omega_{H}=\omega_{0011}$. Use Eq. (23) for the subdominant mode, one obtains,
@@
On the other hand, the GW emission rate of the subdominant mode in Eq. (28a) gives,
@@
where Eq. (43) is used for $M_{\text{v}}^{(0011)}$. Compared to $\dot{E}_{\text{H}}^{(1011)}$, the GW emission rate $\dot{E}^{(1011)}_{\mathrm{GW}}$ has one more power of $M\mu$, but the overall coefficient is larger by a factor of 9. Inserting both terms into Eq. (41c), we see that after $t_{1}$, the mass $M_{\text{v}}^{(1011)}$ decreases exponentially as,
@@
which could be used to estimate the lifetime of the subdominant mode. For the order of magnitude, one could take the $M\mu\to 0$ limit and obtain,
@@
Next, we study the dominant mode in the second phase. Combining Eqs. (37a), (54) and (55), one could obtain,
@@
The GW emission rate $\dot{E}^{(0011)}_{\mathrm{GW}}$ has been calculated in Eq. (26). The ratio of the two rates is,
@@
The vector and scalar superradiance have very different behavior in this phase. Instead of the exponential decay, the scalar dominant mode experiences a second superradiant growth Guo:2022mpr  [^1] Ref. Guo:2022mpr used the single-mode GW emission rate, the counterpart of Eq. (26), for the subdominant mode, while the strongest GW emission comes from the interference term, the counterpart of Eq. (28a). The qualitative behavior does not change after the correction.. This is attributed to the fact that in the scalar case, the GW emission rate is $\sim(M_{1}\mu)^{4}$ smaller, while the suppression of the superradiance rate is only $\sim(M_{1}\mu)^{2}$. As a result, the ratio in Eq. (61) is enhanced by $(M_{1}\mu)^{-2}$ for scalars. The ratio turns out to be larger than 1, causing a second superradiant growth of the scalar dominant mode.
The third phase starts from $t_{2}$ and ends at the time when the $j=2$ modes become important. The BH spin is slightly below $a_{*\text{c}}^{(0011)}$ throughout this phase. The dominant mode with $j=1$ shrinks. Most of its energy is emitted as GW, while a small fraction is returned to the BH. The BH mass is at first $M_{1}$ but then presents a tiny local peak structure. The local peak forms because a vector in the $j=2$ modes takes twice the angular momentum as a vector in the $j=1$ modes. When the energy transfers from the $j=1$ dominant mode to the $j=2$ modes, some energy must be returned to the BH to keep the constant BH spin. We name the time of the local peak as $t_{3}$ and use it as the end of the third phase.
@@ IV.2 The $j=2$ stage
From $t_{3}$ to $t_{1}^{\prime}$, the BH mass drops from $M_{1}$ to,
@@
For the order of magnitude, we could set the logarithm to 10 and replace $M_{1}$ by $M_{0}$, which gives,
@@
The second phase starts at $t_{1}^{\prime}$ and ends at $t_{2}^{\prime}$ when the ${}^{3}P_{2,2}^{1}$ is drained. In this phase, the BH spin is between $a_{*\text{c}}^{(1122)}$ and $a_{*\text{c}}^{(0122)}$. At $t_{2}^{\prime}$, the BH spin equals to $a_{*\text{c}}^{(0122)}$. In this phase, the BH mass is almost unchanged with value $M_{2}$. Following the same steps as in the $j=1$ stage, one could show $M_{\text{v}}^{(1122)}$ decreases exponentially as,
@@
For the order of magnitude, one could take the $M\mu\to 0$ limit and replace $M_{1}$ by $M_{0}$, which gives,
@@
In this phase, the evolution of the ${}^{3}P_{2,2}^{0}$ mode is dominated by the GW emission, the same as the $j=1$ stage. Consequently, its mass can be estimated as,
@@
The third phase starts from $t_{2}^{\prime}$ and ends at the time when the $j=3$ modes become important. The BH spin is slightly below $a_{*\text{c}}^{(0122)}$ throughout this phase. The ${}^{3}P_{2,2}^{0}$ mode shrinks. Its mass is still described by Eq. (71). Most of its energy is emitted as GW, while a small fraction is returned to the BH. The BH mass is at first $M_{2}$ but then presents a tiny local peak structure. The local peak forms because a vector in the $j=3$ modes takes 1.5 times the angular momentum as a vector in the $j=2$ modes. When the energy transfers from the $j=2$ dominant mode to the $j=3$ modes, some energy must be returned to the BH to keep the constant BH spin. We name the time of the local peak as $t_{3}^{\prime}$ and use it as the end of the third phase.
@@ IV.3 Entire Evolution
In this part, we numerically solve the evolution equations in Eqs. (41) and compare the results with the analytic approximations obtained above. The initial BH mass and spin are set as $M_{0}\mu=0.1$ and $a_{*0}=0.99$, respectively. Two $j=1$ modes (${}^{3}S_{1,1}^{0}$ and ${}^{3}S_{1,1}^{1}$) and two $j=2$ modes (${}^{3}P_{2,2}^{0}$ and ${}^{3}P_{2,2}^{1}$) are included in the numerical calculation. The initial mass of the four modes are all set as $10^{-5}M_{0}$. The time dependence of the mass of each mode, the BH mass and spin, as well as the GW flux, are shown in Fig. 8. The comparison of the analytic approximations and the numerical values are further exhibited in Tab. 4.
@@
Quantities $M_{\mathrm{v}}^{(0011)}(t_{1})/M_{0}$ $M_{\mathrm{v}}^{(1011)}(t_{1})/M_{0}$ $\tau_{\mathrm{sr}}^{(0011)}/M_{0}$ $\tau_{\mathrm{life}}^{(1011)}/M_{0}$ $\tau_{\mathrm{GW}}^{(0011)}/M_{0}$ $M_{1}/M_{0}$ $a_{*1}$
@@
Quantities $M_{\mathrm{v}}^{(0122)}(t_{1}^{\prime})/M_{0}$ $M_{\mathrm{v}}^{(1122)}(t_{1}^{\prime})/M_{0}$ $\tau_{\mathrm{sr}}^{(0122)}/M_{0}$ $\tau_{\mathrm{life}}^{(1122)}/M_{0}$ $\tau_{\mathrm{GW}}^{(0122)}/M_{0}$ $M_{2}/M_{0}$ $a_{*2}$
@@
    Subdominant Attractor phase: In the time range $t_{1}<t\leqslant t_{2}$, with $t_{2}=2.23\times 10^{10}M_{0}$, the subdominant ${}^{3}S_{1,1}^{1}$ shrinks exponentially until is drained at $t_{2}$. Numerically, we call a mode is drained when its mass is below $10^{-8}M_{0}$. The BH mass and spin remain roughly constant at $0.933M_{0}$ and $0.359$, respectively. Mathematically, the BH is at the attractor because of the presence of the ${}^{3}S_{1,1}^{1}$ mode. Meanwhile, the dominant ${}^{3}S_{1,1}^{0}$ mode loses $6.435\times 10^{-2}M_{0}$ in mass, and dissipate $6.442\times 10^{-2}M_{0}$ via GW emission. The small discrepancy is due to the extraction of energy from the BH. It confirms the argument that the dominant mode in this phase is governed by the GW emission. In contrast, the subdominant ${}^{3}S_{1,1}^{1}$ mode loses $2.99\times 10^{-5}M_{0}$ in mass, more than 3 times larger than the energy $7.23\times 10^{-6}M_{0}$ emitted as GW. The energy difference is absorbed by the BH. Thus both the absorption and the GW emission are important for the evolution of the subdominant mode in this phase.
@@
Figure 9: The evolutions of the vector condensate mass (panel a), BH mass (panel b), BH spin (panel c), and total GW emission flux (panel d) as functions of time. Two dominant modes (${}^{3}S_{1,1}^{0}$ and ${}^{3}P_{2,2}^{0}$) and two subdominant modes (${}^{3}S_{1,1}^{1}$ and ${}^{3}P_{2,2}^{1}$) are included in the evolution. The initial parameters are $a_{*0}=0.99$ and $M_{0}\mu=0.1$. The initial masses of the $j=1$ modes are chosen as $10^{-5}M_{0}$ and $10^{-1}M_{0}$. Numerical calculation confirms that the initial masses of small-$j$ modes have no impact on the large-$j$ modes, allowing the direct connection between the initial BH parameters and the large-$j$ modes evolution.
@@ V The detection of vector superradiance
Figure 10: The strength and the duration of the GW beat with different BH initial parameters. The initial masses of the vector condensates are $M_{\text{v}0}^{\text{(0011)}}=M_{\text{v}0}^{\text{(1011)}}=10^{-5}M_{0}$ (left panel) and $M_{\text{v}0}^{\text{(0011)}}=M_{\text{v}0}^{\text{(1011)}}=10^{-3}M_{0}$ (right panel). The red dotted line is $a_{*0}=a^{(0011)}_{\mathrm{*c}}$, below which the ${}^{3}S_{1,1}^{0}$ superradiance does not exist. The blue, orange, and pink shaded regions are for $\left(M^{(0011)}_{\mathrm{v,max}}+M^{(1011)}_{\mathrm{v,max}}\right)/M_{1}$ being greater than 0.001, 0.01, 0.05, respectively. The solid contours are calculated with $M^{(1011)}_{\mathrm{v,max}}/M^{(0011)}_{\mathrm{v,max}}$ being the constants labelled above the curves. The dashed contours are for constant $\tau_{\mathrm{life}}^{(1011)}$ labeled beneath the curves, with values in the unit of $M_{0}$. The conversion from $M_{0}$ to the SI units is $M_{\odot}=4.92\times 10^{-6}$ s.

### v2

@@ IV Multi-mode evolution
For the $j=1$ modes, the superradiance rate in Eq. (23) is proportional to $(M\mu)^{7}$, while the GW emission rate is proportional to $(M\mu)^{10}$. For the $j=2$ modes, the superradiance rate is $\sim(M\mu)^{11}$, and the emission rate is $\sim(M\mu)^{14}$. In the $M\mu\ll 1$ limit, the $j=2$ modes become crucial at the time when the $j=1$ modes are almost drained by GW emission. As a result, the evolution of these four modes can be separated into two stages. In the first stage, only the $j=1$ modes are important while the $j=2$ modes can be ignored. Then in the second stage, the $j=1$ modes are almost depleted and one only needs to consider the $j=2$ modes. Such hierarchy also leads to different phases in each stage similar to the single-mode evolution. In fact, we will see that the effect of the subdominant mode is so small that the evolutions of the dominant mode and the BH are almost the same as the single-mode case detailed in Sec. III.
@@ IV.1 The $j=1$ stage
where $\Gamma_{1011}/\Gamma_{0011}\sim\beta_{10}/\beta_{00}=1/8$ with
@@
It means the mass ratio of the subdominant mode to the dominant mode at $t_{1}$ is smaller if the latter is larger in size. This ratio is important for the modulation of the GW emission in Eq. (30), which will be discussed in Sec. V.
@@
During this time, the BH mass and spin drop rapidly to $M_{1}$ and $a_{*\text{c}}^{(1011)}$, respectively. Since the subdominant mode is much smaller in mass than the dominant one, one could make estimates of $M_{1}$ and $M_{\text{v}}^{0011}(t_{1})$ as if only the dominant mode exists. Then the problem reduces to the single-mode evolution explained in Sec. III.1. $M_{\text{v,max}}^{(0011)}$ and $a_{*\text{c}}^{(0011)}$ in Eqs. (42) should be replaced by $M_{\text{v}}^{(0011)}(t_{1})$ and $a_{*\text{c}}^{(1011)}$, respectively. Nonetheless, using $a_{*\text{c}}^{(1011)}$ or $a_{*\text{c}}^{(0011)}$ here leads to an error only at order $\mathcal{O}\left((M\mu)^{8}\right)$. As a result, Eq. (43) can be used to estimate $M_{\text{v}}^{(0011)}(t_{1})$.
With $M_{\text{v}}^{(0011)}(t_{1})$ calculated, one could further obtain $M_{\text{v}}^{(1011)}(t_{1})$ with Eq. (51) and obtain $t_{1}$ with Eq. (50). Apparently, $M_{\text{v}}^{(1011)}(t_{1})$ is the maximum mass of the subdominant mode. The BH mass at time $t_{1}$ is then
$M_{1}$ $\approx M_{0}-{M_{\mathrm{v}}^{(1011)}}(t_{1})-{M_{\mathrm{v}}^{(0011)}}(t_{1}).$ (53)
The second phase starts at $t_{1}$ and ends at $t_{2}$ when the subdominant mode is drained. In this phase, the BH spin is between $a_{*\text{c}}^{(1011)}$ and $a_{*\text{c}}^{(0011)}$. The subdominant mode gradually falls into the BH while the dominant mode still extracts energy out of the horizon. At $t_{2}$, the BH spin equals to $a_{*\text{c}}^{(0011)}$. In this phase, the BH mass is almost unchanged, meaning the balance of the inflow and the outflow at the horizon
@@
Since $M_{\mathrm{v}}^{(0011)}(t)\gg M_{\mathrm{v}}^{(1011)}(t)$, the BH spin in this time range is very close to $a_{*\text{c}}^{(0011)}$. As an approximation, we assume the BH spin is always $a_{*\text{c}}^{(0011)}$ in this time range. Then the superradiance rate $\Gamma_{0011}=0$, indicating the horizon angular speed $\Omega_{H}=\omega_{0011}$. Use Eq. (23) for the subdominant mode, one obtains
@@
On the other hand, the GW emission rate of the subdominant mode in Eq. (28a) gives
@@
where Eq. (43) is used for $M_{\text{v}}^{(0011)}$. Compared to $\dot{E}_{\text{H}}^{(1011)}$, the GW emission rate $\dot{E}^{(1011)}_{\mathrm{GW}}$ has one more power of $M\mu$, but the overall coefficient is larger by a factor of 9. Inserting both terms into Eq. (41c), we see that after $t_{1}$, the mass $M_{\text{v}}^{(1011)}$ decreases exponentially as
@@
which could be used to estimate the lifetime of the subdominant mode. For the order of magnitude, one could take the $M\mu\to 0$ limit and obtain
@@
Next, we study the dominant mode in the second phase. Combining Eqs. (37a), (54) and (55), one could obtain
@@
The GW emission rate $\dot{E}^{(0011)}_{\mathrm{GW}}$ has been calculated in Eq. (26). The ratio of the two rates is
@@
The vector and scalar superradiance have very different behavior in this phase. Instead of the exponential decay, the scalar dominant mode experiences a second superradiant growth Guo:2022mpr . [^1] Ref. Guo:2022mpr used the single-mode GW emission rate, the counterpart of Eq. (26), for the subdominant mode, while the strongest GW emission comes from the interference term, the counterpart of Eq. (28a). The qualitative behavior does not change after the correction. This is attributed to the fact that in the scalar case, the GW emission rate is $\sim(M_{1}\mu)^{4}$ smaller, while the suppression of the superradiance rate is only $\sim(M_{1}\mu)^{2}$. As a result, the ratio in Eq. (61) is enhanced by $(M_{1}\mu)^{-2}$ for scalars. The ratio turns out to be larger than 1, causing a second superradiant growth of the scalar dominant mode.
The third phase starts from $t_{2}$ and ends at the time when the $j=2$ modes become important. The BH spin is slightly below $a_{*\text{c}}^{(0011)}$ throughout this phase. The dominant mode with $j=1$ shrinks. Most of its energy is emitted as GWs, while a small fraction is returned to the BH. The BH mass is at first $M_{1}$ but then presents a tiny local peak structure. The local peak forms because a vector in the $j=2$ modes takes twice the angular momentum as a vector in the $j=1$ modes. When the energy transfers from the $j=1$ dominant mode to the $j=2$ modes, some energy must be returned to the BH to keep the constant BH spin. We name the time of the local peak as $t_{3}$ and use it as the end of the third phase.
@@ IV.2 The $j=2$ stage
From $t_{3}$ to $t_{1}^{\prime}$, the BH mass drops from $M_{1}$ to
@@
For the order of magnitude, we could set the logarithm to 10 and replace $M_{1}$ by $M_{0}$, which gives
@@
The second phase starts at $t_{1}^{\prime}$ and ends at $t_{2}^{\prime}$ when the ${}^{3}P_{2,2}^{1}$ mode is drained. In this phase, the BH spin is between $a_{*\text{c}}^{(1122)}$ and $a_{*\text{c}}^{(0122)}$. At $t_{2}^{\prime}$, the BH spin equals to $a_{*\text{c}}^{(0122)}$. In this phase, the BH mass is almost unchanged with value $M_{2}$. Following the same steps as in the $j=1$ stage, one could show $M_{\text{v}}^{(1122)}$ decreases exponentially as
@@
For the order of magnitude, one could take the $M\mu\to 0$ limit and replace $M_{1}$ by $M_{0}$, which gives
@@
In this phase, the evolution of the ${}^{3}P_{2,2}^{0}$ mode is dominated by the GW emission, the same as the $j=1$ stage. Consequently, its mass can be estimated as
@@
The third phase starts from $t_{2}^{\prime}$ and ends at the time when the $j=3$ modes become important. The BH spin is slightly below $a_{*\text{c}}^{(0122)}$ throughout this phase. The ${}^{3}P_{2,2}^{0}$ mode shrinks. Its mass is still described by Eq. (71). Most of its energy is emitted as GWs, while a small fraction is returned to the BH. The BH mass is at first $M_{2}$ but then presents a tiny local peak structure. The local peak forms because a vector in the $j=3$ modes takes 1.5 times the angular momentum as a vector in the $j=2$ modes. When the energy transfers from the $j=2$ dominant mode to the $j=3$ modes, some energy must be returned to the BH to keep the constant BH spin. We name the time of the local peak as $t_{3}^{\prime}$ and use it as the end of the third phase.
@@ IV.3 Entire evolution
In this part, we numerically solve the evolution equations in Eqs. (41) and compare the results with the analytic approximations obtained above. The initial BH masses and spin are set as $M_{0}\mu=0.1$ and $a_{*0}=0.99$, respectively. Two $j=1$ modes (${}^{3}S_{1,1}^{0}$ and ${}^{3}S_{1,1}^{1}$) and two $j=2$ modes (${}^{3}P_{2,2}^{0}$ and ${}^{3}P_{2,2}^{1}$) are included in the numerical calculation. The initial mass of the four modes are all set as $10^{-5}M_{0}$. The time dependence of the mass of each mode, the BH mass and spin, as well as the GW flux, are shown in Fig. 8. The comparison of the analytic approximations and the numerical values are further exhibited in Table. 4.
@@
Quantities $M_{\mathrm{v}}^{(0011)}(t_{1})/M_{0}$ $M_{\mathrm{v}}^{(1011)}(t_{1})/M_{0}$ $t_{1}/M_{0}$ $\tau_{\mathrm{life}}^{(1011)}/M_{0}$ $\tau_{\mathrm{GW}}^{(0011)}/M_{0}$ $M_{1}/M_{0}$ $a_{*1}$
@@
Quantities $M_{\mathrm{v}}^{(0122)}(t_{1}^{\prime})/M_{0}$ $M_{\mathrm{v}}^{(1122)}(t_{1}^{\prime})/M_{0}$ $t_{1}^{\prime}/M_{0}$ $\tau_{\mathrm{life}}^{(1122)}/M_{0}$ $\tau_{\mathrm{GW}}^{(0122)}/M_{0}$ $M_{2}/M_{0}$ $a_{*2}$
@@
    Subdominant attractor phase: In the time range $t_{1}<t\leqslant t_{2}$, with $t_{2}=2.23\times 10^{10}M_{0}$, the subdominant ${}^{3}S_{1,1}^{1}$ mode shrinks exponentially until is drained at $t_{2}$. Numerically, we call a mode is drained when its mass is below $10^{-8}M_{0}$. The BH mass and spin remain roughly constant at $0.933M_{0}$ and $0.359$, respectively. Mathematically, the BH is at the attractor because of the presence of the ${}^{3}S_{1,1}^{1}$ mode. Meanwhile, the dominant ${}^{3}S_{1,1}^{0}$ mode loses $6.435\times 10^{-2}M_{0}$ in mass, and dissipates $6.442\times 10^{-2}M_{0}$ via GW emission. The small discrepancy is due to the extraction of energy from the BH. It confirms the argument that the dominant mode in this phase is governed by the GW emission. In contrast, the subdominant ${}^{3}S_{1,1}^{1}$ mode loses $2.99\times 10^{-5}M_{0}$ in mass, more than 3 times larger than the energy $7.23\times 10^{-6}M_{0}$ emitted as GW. The energy difference is absorbed by the BH. Thus both the absorption and the GW emission are important for the evolution of the subdominant mode in this phase.
@@
Figure 9: The evolutions of the vector condensate mass (panel a), BH mass (panel b), BH spin (panel c), and total GW emission flux (panel d) as functions of time. Two dominant modes (${}^{3}S_{1,1}^{0}$ and ${}^{3}P_{2,2}^{0}$) and two subdominant modes (${}^{3}S_{1,1}^{1}$ and ${}^{3}P_{2,2}^{1}$) are included in the evolution. The initial parameters are $a_{*0}=0.99$ and $M_{0}\mu=0.1$. The initial masses of the $j=1$ modes are chosen as $10^{-5}M_{0}$ and $10^{-1}M_{0}$. Numerical calculation confirms that the initial masses of small-$j$ modes have no impact on the large-$j$ modes, allowing the direct connection between the initial BH parameters and the large-$j$ mode evolution.
@@ V The detection of vector superradiance
Figure 10: The strength and the duration of the GW beat with different BH initial parameters. The initial masses of the vector condensates are $M_{\text{v}0}^{\text{(0011)}}=M_{\text{v}0}^{\text{(1011)}}=10^{-5}M_{0}$ (panel a) and $M_{\text{v}0}^{\text{(0011)}}=M_{\text{v}0}^{\text{(1011)}}=10^{-3}M_{0}$ (panel b). The red dotted line is $a_{*0}=a^{(0011)}_{\mathrm{*c}}$, below which the ${}^{3}S_{1,1}^{0}$ superradiance does not exist. The blue, orange, and pink shaded regions are for $\left(M^{(0011)}_{\mathrm{v,max}}+M^{(1011)}_{\mathrm{v,max}}\right)/M_{1}$ being greater than 0.001, 0.01, 0.05, respectively. The solid contours are calculated with $M^{(1011)}_{\mathrm{v,max}}/M^{(0011)}_{\mathrm{v,max}}$ being the constants labelled above the curves. The dashed contours are for constant $\tau_{\mathrm{life}}^{(1011)}$ labeled beneath the curves, with values in the unit of $M_{0}$. The conversion from $M_{0}$ to the SI units is $M_{\odot}=4.92\times 10^{-6}$ s.